### 21 Jan 2024 v1

@@ 1 Introduction
In the Engel curve example, the shares of food and leisure are significantly correlated because their sum is less than one, and they should not be treated independently.
@@
Recall that the existing work on endogenous quantile models imposes different monotonicity assumptions for identification.
@@
This restriction implies that the structural function is a multivatriate quantile function that is defined by [6] in the optimal transport literature.
Under other modelling assumptions of [7], we provide conditions for the local identification of the structural function.
@@
For -dimensional potential outcomes, it seems natural that the support size of an IV needed for the identification is greater than two because there are functions to be identified.
@@
The structural function of our model turns out to satisfy a system of functional equations, which we exploit to show the identification.
@@
Technically, this paper is relevant to the optimal transport theory ([29] and [30]) and a partial differential equation (PDE), called the Monge-Ampère equation ([16]).
The structural function of our model is thought of as an optimal-transport-based multivariate quantile function that is investigated in [6], [20], and [17].
@@ 3 IV Multivariate Quantile Model
This restriction can be relaxed, but we keep it for the simplicity.
The condition (2) implies that the IV is independent of potential outcomes conditional on covariates.
The condition (3) is a weak restriction that allows a broad class of assignment rules of treatments.
@@
The condition (4) is called the rank similarity, which is usually assumed in the literature.
@@
Notice that by the condition (1), the observed outcome is written as where
@@
That is, any candidate of structural functions necessarily satisfy (3) in which is replaced by
@@
This is a critical simplification that occurs only in the one-dimensional case in the sense that the set which is intrinsically an infinite dimensional object, is pined down by a scaler
Thanks to this dimension reduction, in order to consider the identification of the structural functions, one can fix a certain quantile level and focus on a finite dimensional system of equations, as [7] do.
@@
To characterize it, one need to know the values of at all the boundary points of which is clearly an infinite dimensional object.
@@
Instead, we consider the identification in an infinite dimensional function space in the next section.
@@ 4 Local Identification of Structural Functions
Notice that this treatment does not affect identification results because the marginal distribution of is identifiable.
@@
Since we are interested in the identification, we assume the correct specification.
@@
The condition (ii) on the eigenvalues that appears in the definition of controls the convexity and smoothness of
@@
The condition (iii) in the definition of excludes pathological candidates of structural functions.
@@
We further impose two regularity conditions.
@@
Suppose that Assumptions 3.1, 4.1, 4.2, 4.3, 4.4 hold.
Then, the structural functions are locally identified in that is, there exists such that if with solves the system (4), then holds.
@@
Assumption 4.4 is a stronger condition than this condition.
@@
Thus, the equation (5) implies
@@
However, it is surprising that the size of the support of the IV, needed for the identification is independent of the dimension of the potential outcomes, as long as the IV is informative enough.
@@ 5 Local Identification for Systems of Equations
Remember that when and are finite dimensional, the local unique solvability of the system of equations is given by the implicit function theorem, which assumes the differentiability and the full rankness of the Jacobian matrix.
As our identification result is an analogue statement for infinite dimensional parameters, we impose assumptions corresponding to these two.
@@
In particular, unlike the Fréchet differentiability, the linearity of is not required, as it may not even be defined on
The next assumption is a infinite dimensional version of the full rankness of the Jacobian matrix.
@@
Our theorem is more general than their result because it covers the local identification problem for systems of equations and because the domain of functionals is not necessarily be a Banach space.
@@
In this section, we outline the proof of Theorem 4.1.
To show Theorem 4.1, we apply Theorem 5.1.
@@
Then, it is clear that which is independent of that and that
@@
Remember the structural function is the optimal transport map from to
If the source and terminal distributions are known, the standard theory ensures the identification of but unfortunately, is not directly observed in our setup.
Still, the support of is identified by Assumption 4.3, which should impose some restrictions on
The following lemma exploits this support condition, which is a key step in our proof.
@@
As the limit function is continuous and almost everywhere on we have up to a subsequence.

### 22 Feb 2024 v2

@@ 1 Introduction
In the Engel curve example, the shares of food and leisure are significantly correlated because there is an obvious substitutability, i.e., more food, less leisure, so that they should not be treated independently.
@@
Existing work on endogenous quantile models imposes different monotonicity assumptions for identification.
@@
Under this restriction, an increase in the th entry of leads to an increase in the th entry of which justifies calling the rank variable vector of
More importantly, a change in a particular entry of can affect other dimensions of both positively and negatively.
This allows the model to capture substitutability/complementarity between different dimensions.
Running the standard IVQR separately does not have this flexibility.
@@
For -dimensional potential outcomes, it may seem natural that the support size of an IV needed for the identification is greater than two because there are functions to be identified.
@@
The structural function of our model turns out to satisfy a system of functional equations.
We exploit this fact to show the identification.
@@
Technically, the cyclical monotonicity assumption on the structural function requires it to be a multivatriate quantile function that is defined and developed by [6], [20], and [17].
Multivariate quantile functions are closely related to the optimal transport theory ([29] and [30]) and a partial differential equation (PDE), called the Monge-Ampère equation ([16]).
Indeed, the structural function of our model turns out to be a solution to a system of Monge-Ampère equations.
The local identification of the structural function can be interpreted as the local uniqueness of the solution to the system, which is one of the central issues in the PDE literature.
We prove the local identification result exploiting the positive definiteness of a differential operator, which is a powerful technique used in the elliptic PDE literature ([18]).
@@ 3 IV Multivariate Quantile Model
This restriction can be relaxed, but we keep it for simplicity.
Condition (2) implies that the IV is independent of potential outcomes conditional on covariates.
Condition (3) is a weak restriction that allows a broad class of assignment rules of treatments.
@@
Condition (4) is called the rank similarity, which is usually assumed in the literature.
@@
Notice that by condition (1), the observed outcome is written as where
@@
That is, any candidate of structural functions necessarily satisfies (4) in which is replaced by
@@
This is a critical simplification that occurs only in the one-dimensional case in the sense that the set which is intrinsically an infinite-dimensional object, is pined down by the scalar
Thanks to this dimension reduction, in order to consider the identification of structural functions, one can fix a certain quantile level and focus on a finite-dimensional system of equations, as [7] do.
@@
To characterize it, one need to know the set of the values of at all the boundary points of which is clearly an infinite-dimensional object.
@@
We consider the identification in an infinite-dimensional function space in the next section.
@@ 4 Local Identification of Structural Functions
Notice that this treatment does not affect identification results because the distributions of observables conditional on are identifiable.
@@
For a compact set in a Euclidean space, we denote the set of continuous functions on by
Also, let be the set of functions on that are -times continuously differentiable on and the derivative can be continuously extended to
@@
Since we are interested in identification, we assume the correct specification.
@@
Condition (ii) on the eigenvalues that appears in the definition of controls the convexity and smoothness of
@@
Condition (iii) in the definition of excludes pathological candidates of structural functions.
@@
We further impose three regularity conditions.
@@
The smoothness of the boundary is rarely a problem in practice, as it allows for kinks at some points.
Combined with Assumption 4.2, Assumption 4.3 implies that is bounded.
@@
Suppose that Assumptions 3.1, 4.1, 4.2, 4.3, 4.4, 4.5 hold.
Then, the structural functions are locally identified in i.e., there exists such that if with solves the system (5), then holds.
@@
Note that Assumption 4.5 requires a positive correlation between and so the labeling of the IV should be switched in favor of the assumption.
@@
Assumption 4.5 is stronger than this condition.
@@
Thus, equation (6) implies
@@
However, it is surprising that the minimum size of the support of the IV, needed for the identification is independent of the dimension of the potential outcomes, as long as the IV is informative enough.
@@ 5 Local Identification for Systems of Equations
Remember that when and are finite-dimensional, the locally unique solvability of the system of equations is given by the implicit function theorem, which assumes the differentiability and the full rankness of the Jacobian matrix.
As our identification result is an analogue statement for infinite-dimensional parameters, we impose assumptions corresponding to these two.
@@
Unlike the Fréchet differentiability, the linearity of is not required.
The next assumption is an infinite-dimensional version of the full rankness of the Jacobian matrix.
@@
Our theorem has a different focus than their result in that it covers the local identification problem for systems of equations and that the domain of functionals is not necessarily be a Banach space.
@@
In this section, we prove Theorem 4.1.
To do so, we apply Theorem 5.1.
@@
As we have which is independent of
Also, it holds that and that
@@
The following lemma, which exploits Assumption 4.4, the identifiability of the support of is a key step in our proof.
Remember that since is continuous, implies
@@
Let be the limit function.
By equation (16), almost everywhere, and it holds everywhere, as is continuous.
Hence, the uniform convergence implies up to a subsequence.